An engine containing 2 kg air as working substance is initially at 1 atm and 27 degree celcius.

The system undergoes isochoric process to a point where pressure of the system is 2

atm. At this point heat is transferred to the air untill the volume doubles. Calulate

the total workdone and the amount of heat transferred to the air.


1
Expert's answer
2021-07-12T02:59:22-0400

W= (nRT/V)ΔV

T = 300K

= 2*8.31*300/(0.5 π 0.12) )*(0.01 π 0.12) J = 279 J